# PE-MA: Parameter-Efficient Co-Evolution of Multi-Agent Systems

### v1

@@ III-B Optimization Objective
where $\ell_{i}$ is the task-specific loss function of agent $a_{i}$ (e.g., cross-entropy or mean squared error), $L_{i}$ is the local optimization objective of agent $a_{i}$ and ${L}$ is the global optimization objective.
@@ IV-B Collaborative Training with Dual-adapter
1:  Initialize: $N$ agents; frozen pretrained model parameters $\boldsymbol{\theta}_{\text{freeze}}$; initial shared adapter $w^{0}$; specific adapters $\{v_{i}^{0}\}_{i=1}^{N}$; local training epochs $\tau$; local datasets $\{\mathbb{D}_{i}\}_{i=1}^{N}$; mixing coefficient $\mu$; learning rates $\eta_{w}$, $\eta_{v}$
@@ V Convergence analysis
$M(t)$ $:=\underset{\text{consensus error of global representation}~{}\mathbf{w}}{\underbrace{\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{w}_{i}(t)-\bar{\mathbf{w}}(t)\|^{2}}}$
@@
The first term in Eq. V measures the average error of the learned global knowledge representation from the perspective of the shared adapter of each agent $a_{i}$. The last two terms represent the performance of our model. We use different learning rates $\eta_{w}$ and $\eta_{v}$ for the shared adapter and the personalized adapter. We are inspired by finite-time analysis of two-timescale stochastic approximation [27] to consider the global loss of weighted partial gradients.
@@
The term $\frac{4L(\bar{\mathbf{w}}(0),\!\{\mathbf{v}_{i}(0)\}_{i=1}^{N})\!-\!4L({\mathbf{w}^{*}},\!\{\mathbf{v}_{i}^{*}\}_{i=1}^{N})}{K\eta_{w}}$ vanishes at a rate proportional to $1/K$, reflecting the initial performance gap between the algorithm and the optimal solution, which gradually diminishes with training. This convergence behavior aligns with that observed in traditional decentralized learning frameworks [28; 29]. In addition, there exists a persistent error term that does not decrease as $K$ increases, representing the long-term “steady-state error” of the algorithm. These errors mainly arise from two sources: (i) variance due to stochastic gradients and model heterogeneity, represented by $\sigma_{1},\sigma_{2}$ and $\varsigma$; and (ii) the discrepancy between each agent’s shared adapter and the average shared adapter, measured by $\|\mathbf{w}_{i}(k)-\bar{\mathbf{w}}(k)\|^{2}$. Notably, the term $\frac{12\eta_{v}^{3}L^{2}\tau}{\eta_{w}}(\tau-1)(6\tau-1)\sigma_{1}^{2}+\frac{2\eta_{v}^{2}\tau L}{\eta_{w}}\sigma_{2}^{2}$ originates from the accumulation of local update errors during the first step. This indicates that each round of local updates progressively introduces errors, and that these errors grow at a higher-order rate with respect to the number of local updates $\tau$.
@@ VI-B Experimental Setup
For experimental comparison, we consider several representative methods, including standard full-model training and personalized collaborative learning strategies based on FedSim [33] and FedAlt [34], applied to three different levels: input layer, outer layer, and adapter layer. During training, the backbone network parameters remain frozen, and only the adapter parameters, encoding head, or fully connected layer parameters are updated depending on the specific method used. The initial learning rate is set to $10^{-2}$, and it becomes $10^{-1}$ of the previous one every three epochs, and the batch size is 64. We implement all algorithms in PyTorch [35] on Python 3 with ten NVIDIA RTX 3090ti GPUs.
@@ Appendix A Convergence
[…]
+$\quad$+ $\frac{L}{2}\|\mathbf{w}-\mathbf{w}^{\prime}\|^{2}+\nabla_{\mathbf{v}}L_{i}(\mathbf{w}^{\prime},\mathbf{v}^{\prime})^{\intercal}(\mathbf{v}-\mathbf{v}^{\prime})+\frac{L}{2}\|\mathbf{v}-\mathbf{v}^{\prime}\|^{2}.$
@@
=$\quad$= $L_{i}(\mathbf{w},\mathbf{v})-L_{i}(\mathbf{w}^{\prime},\mathbf{v})+L_{i}(\mathbf{w}^{\prime},\mathbf{v})-L_{i}(\mathbf{w}^{\prime},\mathbf{v}^{\prime})$
[…]
@@
[…]
@@
=$\quad$= $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[L_{i}(\mathbf{w}(t+1),\mathbf{v}_{i}(t+1))]\allowdisplaybreaks$
[…]
@@
=$\overset{(a_{1})}{=}\quad$ $-\eta_{w}\mathbb{E}\Big{\langle}\nabla_{\mathbf{w}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N}),$
@@
=$\overset{(a_{2})}{=}\quad$ $\frac{-\eta_{w}}{2}\Bigg{(}\left\|\nabla_{\mathbf{w}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N})\right\|^{2}$ (17)
@@
=$\quad$= $\frac{\eta_{w}}{2N^{2}}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{i=1}^{N}(\nabla_{\mathbf{w}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t+1))$
@@
=$\overset{(b_{1})}{=}\quad$ $\frac{\eta_{w}^{2}L}{2N^{2}}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{i=1}^{N}g_{\mathbf{w}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t+1))$
@@
=$\quad$= $\frac{-\eta_{v}\tau}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))\right\|^{2}\right]$
@@
=$\overset{(c_{1})}{=}\quad$ $\frac{-\eta_{v}\tau}{N}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))Bigg\|^{2}\Bigg{]}$
@@
=$\quad$= $\frac{-\eta_{v}\tau}{2N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))\right\|^{2}\right]$
@@
=$\quad$= $\frac{\eta_{v}}{2N\tau}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))$
@@
[…]
@@
[…]
@@
=$\quad$= $\mathbb{E}\Bigg{[}\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)-\eta_{v}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s-1))\|^{2}\Bigg{]}\allowdisplaybreaks$
=$\quad$= $\mathbb{E}\Big{[}\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)-\eta_{v}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s-1))$
@@
[…]
@@
[…]
@@
[…]
@@
=$\quad$= $\frac{\eta_{v}^{2}L}{2N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))\right\|^{2}\right]$
[…]
@@
[…]
@@
[…]
@@
=$\quad$= $\mathbb{E}\Bigg{\|}\frac{1}{N}\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)-\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)P^{k}(i,j)$
@@
=$\quad$= $\mathbb{E}\Bigg{\|}\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)\Bigg{(}\frac{1}{N}-P^{k}(i,j)\Bigg{)}$
@@
=$\overset{(e_{3})}{=}\quad$ $\mathbb{E}\Bigg{\|}\eta_{w}C\sum\limits_{r=0}^{k-1}q^{(t-r)}\sum\limits_{j=1}^{N}g_{\mathbf{w}}(\mathbf{w}_{j}(r),\mathbf{v}_{j}(r+1))\Bigg{\|}^{2}$
=$\quad$= $\mathbb{E}\Bigg{\|}\eta_{w}C\sum\limits_{r=0}^{k-1}q^{(t-r)}\sum\limits_{j=1}^{N}(g_{\mathbf{w}}(\mathbf{w}_{j}(r),\mathbf{v}_{j}(r+1))$
@@
[…]
@@
[…]
@@
[…]
@@
+$\underset{(2):(1)+(2)\leq 0}{\underbrace{\quad+\frac{18\eta_{v}^{3}L^{2}(\tau^{3}-\tau^{2})}{N}\sum_{i=1}^{N}\mathbb{E}\left\|g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t))\right\|^{2}}}\allowdisplaybreaks$
@@
[…]
@@
[…]
@@
[…]
@@
where the second inequality holds due to the initialization such that $\sum_{i=1}^{N}\mathbb{E}\|\mathbf{w}_{i}(0)-\bar{\mathbf{w}}(t)\|^{2}=0.$ Summing the recursion in (A) from round 00 to round $K-1$ yields
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
=$\quad$= $\frac{1}{K}\sum\limits_{k=0}^{K-1}\frac{\eta_{v}\tau}{\eta_{w}}\mathbb{E}\left\|\nabla_{\mathbf{v}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})\right\|^{2}$
@@
[…]
@@
[…]

### v2

@@ III-B Optimization Objective
where $\ell_{i}$ is the task-specific loss function of agent $a_{i}$ (e.g., cross-entropy or mean squared error), $L_{i}$ is the local optimization objective of agent $a_{i}$ and L{L} is the global optimization objective.
@@ IV-B Collaborative Training with Dual-adapter
1:Initialize: NN agents; frozen pretrained model parameters $\boldsymbol{\theta}_{\text{freeze}}$; initial shared adapter $w^{0}$; specific adapters $\{v_{i}^{0}\}_{i=1}^{N}$; local training epochs $\tau$; local datasets $\{\mathbb{D}_{i}\}_{i=1}^{N}$; mixing coefficient $\mu$; learning rates $\eta_{w}$, $\eta_{v}$
@@ V Convergence analysis
$M(t)$ $=\underset{\text{consensus error of global representation}~\mathbf{w}}{\underbrace{\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{w}_{i}(t)-\bar{\mathbf{w}}(t)\|^{2}}}$
@@
The first term in Eq. V measures the average error of the learned global knowledge representation from the perspective of the shared adapter of each agent $a_{i}$. The last two terms represent the performance of our model. We use different learning rates $\eta_{w}$ and $\eta_{v}$ for the shared adapter and the personalized adapter. We are inspired by finite-time analysis of two-timescale stochastic approximation [27] to consider the global loss of weighted partial gradients [28].
@@
The term $\frac{4L(\bar{\mathbf{w}}(0),\!\{\mathbf{v}_{i}(0)\}_{i=1}^{N})\!-\!4L({\mathbf{w}^{*}},\!\{\mathbf{v}_{i}^{*}\}_{i=1}^{N})}{K\eta_{w}}$ vanishes at a rate proportional to $1/K$, reflecting the initial performance gap between the algorithm and the optimal solution, which gradually diminishes with training. This convergence behavior aligns with that observed in traditional decentralized learning frameworks [29; 30]. In addition, there exists a persistent error term that does not decrease as $K$ increases, representing the long-term “steady-state error” of the algorithm. These errors mainly arise from two sources: (i) variance due to stochastic gradients and model heterogeneity, represented by $\sigma_{1},\sigma_{2}$ and $\varsigma$; and (ii) the discrepancy between each agent’s shared adapter and the average shared adapter, measured by $\|\mathbf{w}_{i}(k)-\bar{\mathbf{w}}(k)\|^{2}$. Notably, the term $\frac{12\eta_{v}^{3}L^{2}\tau}{\eta_{w}}(\tau-1)(6\tau-1)\sigma_{1}^{2}+\frac{2\eta_{v}^{2}\tau L}{\eta_{w}}\sigma_{2}^{2}$ originates from the accumulation of local update errors during the first step. This indicates that each round of local updates progressively introduces errors, and that these errors grow at a higher-order rate with respect to the number of local updates $\tau$.
@@ VI-B Experimental Setup
For experimental comparison, we consider several representative methods, including standard full-model training and personalized collaborative learning strategies based on FedSim [34] and FedAlt [35], applied to three different levels: input layer, outer layer, and adapter layer. During training, the backbone network parameters remain frozen, and only the adapter parameters, encoding head, or fully connected layer parameters are updated depending on the specific method used. The initial learning rate is set to $10^{-2}$, and it becomes $10^{-1}$ of the previous one every three epochs, and the batch size is 64. We implement all algorithms in PyTorch [36] on Python 3 with ten NVIDIA RTX 3090 GPUs.
@@ Appendix A Convergence
$\leq\quad$ $\nabla_{\mathbf{w}}L_{i}(\mathbf{w}^{\prime},\mathbf{v})^{\intercal}(\mathbf{w}-\mathbf{w}^{\prime})$
$+\quad$ $\frac{L}{2}\|\mathbf{w}-\mathbf{w}^{\prime}\|^{2}+\nabla_{\mathbf{v}}L_{i}(\mathbf{w}^{\prime},\mathbf{v}^{\prime})^{\intercal}(\mathbf{v}-\mathbf{v}^{\prime})+\frac{L}{2}\|\mathbf{v}-\mathbf{v}^{\prime}\|^{2}.$
@@
$=\quad$ $L_{i}(\mathbf{w},\mathbf{v})-L_{i}(\mathbf{w}^{\prime},\mathbf{v})+L_{i}(\mathbf{w}^{\prime},\mathbf{v})-L_{i}(\mathbf{w}^{\prime},\mathbf{v}^{\prime})$
$\leq\quad$ $\nabla_{\mathbf{w}}L_{i}(\mathbf{w}^{\prime},\mathbf{v})^{\intercal}(\mathbf{w}-\mathbf{w}^{\prime})+\frac{L}{2}\|\mathbf{w}-\mathbf{w}^{\prime}\|^{2}$
@@
$\leq\quad$ $\underset{C_{1}}{\underbrace{\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\Big{\langle}\nabla_{w}L_{i}(\mathbf{w}(t),\mathbf{v}_{i}(t+1)),\mathbf{w}(t+1)-\mathbf{w}(t)\Big{\rangle}}}$
@@
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[L_{i}(\mathbf{w}(t+1),\mathbf{v}_{i}(t+1))]\allowdisplaybreaks$
$\leq\quad$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[\nabla_{w}L_{i}(\mathbf{w}(t),\mathbf{v}_{i}(t+1))^{\intercal}(\mathbf{w}(t+1)-\mathbf{w}(t))]$
@@
$\overset{(a_{1})}{=}\quad$ $-\eta_{w}\mathbb{E}\Big{\langle}\nabla_{\mathbf{w}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N}),$
@@
$\overset{(a_{2})}{=}\quad$ $\frac{-\eta_{w}}{2}\Bigg{(}\left\|\nabla_{\mathbf{w}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N})\right\|^{2}$ (17)
@@
$=\quad$ $\frac{\eta_{w}}{2N^{2}}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{i=1}^{N}(\nabla_{\mathbf{w}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t+1))$
@@
$\overset{(b_{1})}{=}\quad$ $\frac{\eta_{w}^{2}L}{2N^{2}}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{i=1}^{N}g_{\mathbf{w}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t+1))$
@@
$=\quad$ $\frac{-\eta_{v}\tau}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))\right\|^{2}\right]$
@@
$\overset{(c_{1})}{=}\quad$ $\frac{-\eta_{v}\tau}{N}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))Bigg\|^{2}\Bigg{]}$
@@
$=\quad$ $\frac{-\eta_{v}\tau}{2N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\nabla_{\mathbf{v}}L_{i}(\bar{\mathbf{w}}(t),\mathbf{v}_{i}(t))\right\|^{2}\right]$
@@
$=\quad$ $\frac{\eta_{v}}{2N\tau}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))$
@@
$\leq\quad$ $\frac{\eta_{v}}{N\tau}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{s=0}^{\tau-1}\nabla_{\mathbf{v}}L_{i}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))$
@@
$\leq\quad$ $\frac{\eta_{v}\tau L^{2}}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\mathbf{w}_{i}(t)-\bar{\mathbf{w}}(t)\right\|^{2}\right]$
@@
$=\quad$ $\mathbb{E}\Bigg{[}\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)-\eta_{v}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s-1))\|^{2}\Bigg{]}\allowdisplaybreaks$
$=\quad$ $\mathbb{E}\Big{[}\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)-\eta_{v}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s-1))$
@@
$\leq\quad$ $\left(1+\frac{1}{2\tau-1}\right)\mathbb{E}\left[\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)\|^{2}\right]$
@@
$\leq\quad$ $\Bigg{(}1+\frac{1}{2\tau-1}\Bigg{)}\mathbb{E}\left[\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)\|^{2}\right]+\eta_{v}^{2}\sigma^{2}$
@@
$\leq\quad$ $\left(1+\frac{1}{2\tau-1}+6\eta_{v}^{2}\tau L^{2}\right)\mathbb{E}\left[\|\mathbf{v}_{i}(t,s-1)-\mathbf{v}_{i}(t)\|^{2}\right]$
@@
$=\quad$ $\frac{\eta_{v}^{2}L}{2N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))\right\|^{2}\right]$
${=}\quad$ $\frac{\eta_{v}^{2}L}{2N}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\Bigg{\|}\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))$
@@
${\leq}\quad$ $\frac{\eta_{v}^{2}\tau L}{2}\sigma_{2}^{2}+\frac{\eta_{v}^{2}L}{2N}\sum_{i=1}^{N}\left\|\mathbb{E}\sum_{s=0}^{\tau-1}g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t,s))\right\|^{2}.$ (28)
@@
${\leq}\quad$ ${\frac{\eta_{w}L^{2}}{2N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\bar{\mathbf{w}}(t)-\mathbf{w}_{i}(t)\right\|^{2}\right]}$
@@
$=\quad$ $\mathbb{E}\Bigg{\|}\frac{1}{N}\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)-\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)P^{k}(i,j)$
@@
$=\quad$ $\mathbb{E}\Bigg{\|}\sum\limits_{j=1}^{N}\mathbf{w}_{j}(0)\Bigg{(}\frac{1}{N}-P^{k}(i,j)\Bigg{)}$
@@
$\overset{(e_{3})}{=}\quad$ $\mathbb{E}\Bigg{\|}\eta_{w}C\sum\limits_{r=0}^{k-1}q^{(t-r)}\sum\limits_{j=1}^{N}g_{\mathbf{w}}(\mathbf{w}_{j}(r),\mathbf{v}_{j}(r+1))\Bigg{\|}^{2}$
$=\quad$ $\mathbb{E}\Bigg{\|}\eta_{w}C\sum\limits_{r=0}^{k-1}q^{(t-r)}\sum\limits_{j=1}^{N}(g_{\mathbf{w}}(\mathbf{w}_{j}(r),\mathbf{v}_{j}(r+1))$
@@
$\leq\quad$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|\nabla_{\mathbf{w}}L_{i}(\mathbf{w}_{i}(t-1),\mathbf{v}_{i}(t))$
@@
$\leq\quad$ $\frac{3L^{2}}{N}\sum_{i=1}^{N}\mathbb{E}\|\mathbf{w}_{i}(t)-\bar{\mathbf{w}}(t)\|^{2}+3\varsigma^{2}$
@@
$\leq\quad$ $\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]-\mathbb{E}[L(\bar{\mathbf{w}}(t+1),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N})]$
@@
$\underset{(2):(1)+(2)\leq 0}{\underbrace{\quad+\frac{18\eta_{v}^{3}L^{2}(\tau^{3}-\tau^{2})}{N}\sum_{i=1}^{N}\mathbb{E}\left\|g_{\mathbf{v}}(\mathbf{w}_{i}(t),\mathbf{v}_{i}(t))\right\|^{2}}}\allowdisplaybreaks$
@@
$\leq\quad$ $\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]-\mathbb{E}[L(\bar{\mathbf{w}}(t+1),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N})]$
@@
$\leq\quad$ $\Bigg{(}\frac{18N^{2}\eta_{w}^{2}C^{2}L^{2}}{(1-q)^{2}}\sum_{k=0}^{K-1}\sum_{i=1}^{N}\mathbb{E}\|\mathbf{w}_{i}(t)-\bar{\mathbf{w}}(t)\|^{2}\allowdisplaybreaks$
@@
$\leq\quad$ $\frac{18\eta_{w}^{2}C^{2}N^{3}}{(1-q)^{2}-18N^{2}\eta_{w}^{2}C^{2}L^{2}}\sum_{k=0}^{K-1}\mathbb{E}\|\nabla_{\mathbf{w}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t+1)\}_{i=1}^{N})\|^{2}$
@@
where the second inequality holds due to the initialization such that $\sum_{i=1}^{N}\mathbb{E}\|\mathbf{w}_{i}(0)-\bar{\mathbf{w}}(t)\|^{2}=0.$ Summing the recursion in (A) from round 0 to round K1K-1 yields
@@
$\leq\quad$ $\mathbb{E}[L(\bar{\mathbf{w}}(0),\{\mathbf{v}_{i}(0)\}_{i=1}^{N})]-\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]$
@@
$\leq\quad$ $\mathbb{E}[L(\bar{\mathbf{w}}(0),\{\mathbf{v}_{i}(0)\}_{i=1}^{N})]-\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]\allowdisplaybreaks$
@@
$\leq\quad$ $\sum\limits_{t=0}^{K-1}\Bigg{(}\frac{\eta_{v}\tau}{2}\left\|\nabla_{\mathbf{v}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})\right\|^{2}$
@@
$\leq\quad$ $\mathbb{E}[L(\bar{\mathbf{w}}(0),\{\mathbf{v}_{i}(0)\}_{i=1}^{N})]-\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]$
@@
$=\quad$ $\frac{1}{K}\sum\limits_{k=0}^{K-1}\frac{\eta_{v}\tau}{\eta_{w}}\mathbb{E}\left\|\nabla_{\mathbf{v}}L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})\right\|^{2}$
@@
$\leq\quad$ $\frac{4\mathbb{E}[L(\bar{\mathbf{w}}(0),\{\mathbf{v}_{i}(0)\}_{i=1}^{N})]-4\mathbb{E}[L(\bar{\mathbf{w}}(t),\{\mathbf{v}_{i}(t)\}_{i=1}^{N})]}{K\eta_{w}}$
@@
$\leq\quad$ $\frac{4f\left(\bar{\mathbf{w}}(0),\{\mathbf{v}_{i}(0)\}_{i=1}^{N}\right)-4f\left({\mathbf{w}}^{*},\{\mathbf{v}_{i}^{*}\}_{i=1}^{N}\right)}{K\eta_{w}}$